a dairyman has milk with 5% butterfat and cream with 20% butterfat.How much of each ingredient must be mixed to make 30 gallons with 6% butterfat?
Answer by stanbon(75887)

a dairyman has milk with 5% butterfat and cream with 20% butterfat.How much of each ingredient must be mixed to make 30 gallons with 6% butterfat?
-----
Equation:
fat + fat = fat
0.05x + 0.20(30-x) = 0.06*30
---
Multiply thru by 100 to get:
5x + 20*30 - 20x = 6*30
-15x = -14*30
---
x = 28 gallons (amt of 5% milk needed in the mixture
30-x = 2 gallons (amt of 20% milk needed in the mixture.
==========
Cheers,
Stan H.